Westerville Company reported the following results from last year’s operations:

Sales $ 2,300,000
Variable expenses 670,000
Contribution margin 1,630,000
Fixed expenses 1,170,000
Net operating income $ 460,000
Average operating assets $ 1,437,500

At the beginning of this year, the company has a $287,500 investment opportunity with the following cost and revenue characteristics:

Sales $ 460,000
Contribution margin ratio 50 % of sales
Fixed expenses $ 161,000

The company’s minimum required rate of return is 15%.

a.What is the residual income of this year’s investment opportunity?

b.If the company pursues the investment opportunity and otherwise performs the same as last year, what residual income will it earn this year?
Answer #1
  1. Computation of residual income

Minimum rate of return = 15%

Average operating assets = $ 287,500

Target Return = $ 287,500 * 15%

                           = $ 43,125

Actual Income = $ 69,000

Residual Income = 69,000 – 43,125

                                = $ 25,875

  1. If the company pursue the investment opportunity, this year’s residual income will be:

Minimum rate of return = 15%

Average operating assets = $ 1,437,500 + 287,500 = $ 1,725,000

Target Return = 1,725,000 * 15%

   = 258,750

Actual Income = $ 460,000 + 69,000 = $ 529,000

Residual Income = $ 529,000 – 258,750

                                = $ 270,250